find the equation of the circle (standard form)
the circle has a diameter with endpoints whose coordinates are (7, -2) and (-3, 5).
Please show work thank you!

Answer by lwsshak3(11628)

find the equation of the circle (standard form)
the circle has a diameter with endpoints whose coordinates are (7, -2) and (-3, 5).
***
Standard form of equation for a circle: (x-h)^2+(y-k)^2=r^2, (h,k)=coordinates of center, r=radius
..
center=midpoint of diameter:(7-3)/2,(-2+5)/2=(2,1.5)
diameter=√((7+3)^2+(-2-5)^2)=√(100+49)=√149 (distance formula)
r=√149/2
r^2=149/4
..
equation: (x-2)^2+(y-1.5)^2=149/4